# Generalized Spectral Clustering of Low-Inertia Power Networks

### v1

@@ II-C Linearized Dynamics
The damping terms $D_{i}>0$ can be represented as the node weights on the dynamic graph. Since $D$ is positive definite ($D\succ 0$) and invertible, the linearized angle dynamics of the non-uniform Kuramoto oscillator around a synchronous solution is given by
@@
where $J$ is the system dynamics matrix. By a change of variable, we can write the linearized system dynamics in terms of a symmetric matrix as
$D^{1/2}D^{1/2}\dot{\Delta\delta}=-\tilde{L}D^{-1/2}D^{1/2}\Delta\delta$
so that the angle dynamics in this coordinate system is given by
$D^{1/2}\dot{\Delta\delta}=-(D^{-1/2}\tilde{L}D^{-1/2})D^{1/2}\Delta\delta.$
The system dynamics matrix $J$ is similar to the symmetric matrix $D^{-1/2}\tilde{L}D^{-1/2}$ via the transformation matrix $P=D^{-1/2}$, hence they share the same real eigenvalues and if $u$ is an eigenvector of $D^{-1/2}\tilde{L}D^{-1/2}$, then $v=Pu=D^{-1/2}u$ is the corresponding eigenvector of $J$. We remark here that the eigenvectors of $J$ are $D$-orthogonal, and for a connected network with voltage angles $\delta\in\overline{\Delta}$, the zero eigenvalue of $J$ has an algebraic multiplicity of $1$ and the corresponding eigenvector is simple.
The eigenvalues and vectors of $J$ are also negative of the eigenvalues and the eigenvectors of the generalized eigenvalue problem of the pair $(\tilde{L},D)$, respectively. The remainder of the paper adopts this framework, which facilitates decoupling of the influence of network connectivity from that of the effect of time constants of the oscillators on the state dynamics. Using the Rayleigh-Ritz variational characterization of the eigenvalues of the pair $(\tilde{L},D)$, it is easy to see that $\lambda_{i}(J)\leq 0$ for all $i$ when $L\succeq 0$ and $D\succ 0$, with $0=\lambda_{1}(J)>\lambda_{2}(J)\geq\cdots\geq\lambda_{n}(J)$ when the graph $\tilde{\mathcal{G}}$ is connected, and $\delta=\delta^{*}$ is locally exponentially stable.
For a network of oscillators with uniform damping (i.e., $D=\alpha I$ for some $\alpha>0$) the $i$th eigenvalue of the system dynamics matrix $J$ is given by $\lambda_{i}(J)=-\lambda_{i}(\tilde{L})/\alpha$, while the eigenvectors corresponding to each eigenvalue are the same as those of $\tilde{L}$ since $D^{-1}$ and $\tilde{L}$ are jointly diagonalizable by the eigenvectors of the Laplacian matrix $\tilde{L}$. Hence, the community structure of the network can be solely determined from the eigenvectors of the network Laplacian matrix $\tilde{L}$. In the general case of a network with non-uniform time constants, matrices $\tilde{L}$ and $D$ play a crucial role in determining the evolution of the states variables and the community structure of the network.
In Section III, we show that the slowest eigenvectors (also referred to as the inter-area modes) of the linearized system dynamics matrix $J$ results in a certain decomposition of the network into dynamically coherent groups. The resulting clusters and the broader framework of spectral clustering are described in the following sections.
III Clustering
III-A The Optimization Problem
@@
is the boundary of the set $\mathcal{S}_{i}$ on the graph $\tilde{G}$, that is, the sum of the edge weights leaving $\mathcal{S}_{i}$ or the total dynamic coupling between $\mathcal{S}_{i}$ and $\mathcal{S}_{i}^{c}$,
@@
is the total damping in the subgraph induced by the set $\mathcal{S}_{i}$. Notice when $D_{i}=1$ for all $i$, $D=I_{n}$, and this term is the cardinality of the set $\mathcal{S}_{i}$ denoted by $|\mathcal{S}_{i}|$.
@@
The optimization problem (5) is an integer program known to be computationally challenging for large scale networks [26], with the size of the feasible set, $k^{n}$, growing exponentially with the size of the network $n$. For any choice of $k$, the solution to (5) corresponds to finding $k$ characteristic vectors in $\{1,0\}^{n}$ with disjoint support [11] that minimizes $\rho(k)$.
@@
$R\left(\tilde{L},D;x_{i}\right)=\frac{x_{i}^{\intercal}\tilde{L}x_{i}}{x_{i}^{\intercal}Dx_{i}}.$
The Rayleigh quotients are well defined for $D\succ 0$ and non-zero $x_{i}$; and the relaxation of (5) is given by
@@
The $k$ vectors minimizing (6) are precisely the generalized eigenvectors corresponding to the $k$ smallest eigenvalues of the pair $(\tilde{L},D)$ [11] with an optimal value of $\psi^{*}(k)=\lambda_{k}(\tilde{L},D)$ – this is a consequence of the generalized Courant-Fischer minimax theorem for the matrix pair $(\tilde{L},D)$ when $\tilde{L}\succeq 0$ and $D\succ 0$ [3, 43].
For a given value of $k$, a power network with a small value of $\psi^{*}(k)$ – the $k$th eigenvalue of the linearized dynamics – implies that the network admits to $k$ loosely coupled partitions with balanced total damping. With $\psi^{*}(k)=0$ if and only if the network has at least $k$ connected components ($k$ islands), which means that the algebraic multiplicity of the zero eigenvalue of the Laplacian matrix of the dynamic graph $\tilde{L}$ is at least $k$. The eigenvectors of the linearized dynamics (4) solves the relaxation of the clustering problem 5.
III-B Generalized Spectral Embedding
Graphs are combinatorial objects whose nodes do not have intrinsic coordinates and typically do not lie in an ambient space [4]. Graph embedding is a process of mapping the nodes of a graph to a low dimensional geometric space that preserves the properties of the network. This geometric space, which is generally Euclidean, is such that the distances between coordinates encode the relationships between the nodes in the network. In this section, we use the generalized eigenvectors of the pair $(\tilde{L},D)$, which encodes properties of the linearized dynamics introduced in Section III-A, to construct such an embedding. We then describe a method for obtaining $k$ dynamically coherent clusters (zones) of a power network based on this graph embedding.
@@ III-B Generalized Spectral Embedding
where $\lambda_{i}=\lambda_{i}(\tilde{L},D)\in\mathbb{R}$ and $x_{i}\in\mathbb{R}^{n}$ are the $i$th eigenvalue and right eigenvector of the generalized eigenvalue problem, respectively. The generalized eigenvalue problem can be thought of as an ordinary eigenvalue problem over a vector space equipped with a $D$-inner product.
@@
4:Compute $X\in\mathbb{R}^{n\times k}$, the first $k$ generalized eigenvectors of the pair $(\tilde{L},D)$
@@
For a power network, the number of control zones (the value of $k$) can be determined by the requirements of the underlying control scheme or pre-specified by the system operator. Ideally, $k$ should be such that the first $k$ generalized eigenvalues $\lambda_{1}(\tilde{L},D),\ldots,\lambda_{k}(\tilde{L},D)$ are small while $\lambda_{k+1}(\tilde{L},D)$ is relatively large. In the absence of a prespecified value of $k$, a common heuristic for choosing $k$ is to choose $k$ with the largest relative spectral gap, that is
@@
There are several justification for this choice of $k$. First, a large relative spectral gap between $\lambda_{k}$ and $\lambda_{k+1}$ ensures that the community structure of the network (the number of clusters) is robust across a wide range of steady-state operating points. Appendix B provides an explicit bound on the worst-case change in the eigenvalues of the linearized oscillator model in terms of the spectral radius of the perturbation to the Laplacian of the dynamic graph and a measure of the definiteness of the pair $(\tilde{L},D)$. This choice of $k$ ensures that the worst of the $k$ clusters, as evaluated using $\phi(\cdot)$, returned by the method is good. When the damping coefficients of the oscillators are uniform (i.e. $D=\alpha I$, for some $\alpha>0$) the relative spectral gap does not depend on the value of $\alpha$.
The embedding of the network constructed using Algorithm 1 is unique when the $k$ smallest eigenvalues of the matrix pair $(\tilde{L},D)$ are distinct, since the eigenspaces corresponding to distinct eigenvalues are disjoint subspaces, the $k$ eigenvectors computed at step 4 of Algorithm 1 are therefore unique up to scaling. We note that the dynamic graph $\tilde{G}$ can be constructed solely from the network parameters and measurements without the need to solve an AC optimal power flow problem. This can significantly reduce the computational complexity of the algorithm to that of computing the eigenvectors of the generalized eigenvalue problem – the most computationally expensive step – which can be computed exactly in polynomial time $\mathcal{O}(n^{3})$ using QZ factorization, or approximately using the inverse power method which converges to a sufficiently good solution in $\mathcal{O}(\log(n))$[11].
@@ IV Results
Figure 1: 22-D Embedding of the IEEE 30-bus test network using the (a) eigenvectors of $\tilde{L}$, (b) the generalized eigenvectors of $(\tilde{L},D)$. The two clusters identified by running $k$-means on the respective embeddings of the network are highlighted. The total edge weights cut by the spectral and generalized spectral clustering solutions on $\tilde{\mathcal{G}}$, are 5.04 and 12.55, respectively, and the corresponding total damping in each cluster is 157.89 and 30.01 for spectral clustering, and 91.33 and 96.57 for generalized spectral clustering.
@@ V Numerical Validation
For each $i\in\mathcal{V}$, we perturb $\omega_{i}$ and determine the effect of this disturbance on the network by measuring the pair-wise coherence between the angular frequency trajectories $\dot{\delta}_{i}(t)$ and $\dot{\delta}_{j}(t)$ for all $j$ in the synchronously rotating reference frame of frequency $\omega_{\text{sync}}$. The measure of coherence is the extension of the cosine similarity (i.e., correlation) from finite-dimensional vectors to functions in $\mathcal{L}_{2}$, see Appendix D for details.
@@
V-A Robustness of Clustering Solution
@@
For each load bus (node) in the network, we generate $1,000$ realizations of uniformly distributed load demands (the natural frequencies of these oscillators) as follows
@@
where $\omega_{i}^{0}$ is the nominal demand (in MW) at node $i$, $\zeta_{i}\sim\mathcal{N}(0,\sigma^{2})$ with variance $\sigma^{2}=5$ MW. We then compute the corresponding net-injections of the generator buses, and the voltage magnitudes and angles across the network by solving the ACOPF problem. Randomization in the load demands across the network induces a structured randomization in the realized steady states of the network, ensuring the operational feasibility of the resulting states. This is consistent with typical power system operations, as randomizing the natural frequencies of all oscillators does not guarantee a feasible or economic operating point and may result in unrealizable states of the network.
Figure 8: Distribution of the largest relative spectral gap for $1,000$ random steady-states of the network, with $\mu=3.1218$ and $\sigma^{2}=0.0213$.
Across all considered operating points, the largest relative spectral gap was consistently between the $5$th and $6$th generalized eigenvalues. Fig. 8 shows the distribution of the largest relative spectral gap of the linearized dynamics over the realized states. The distribution has a mean and variance of $3.1218$ and variance of $0.0213$ (relative variance of $0.007$), confirming that choice of $k=5$ clusters for this network is robust to significant variability in system operating conditions.
@@
The robustness of the resulting clusters ensures that the identified clusters remain stable under changes in the network’s steady-state operating points, precluding the need to re-solve the clustering problem for slight changes in the network’s loading condition. This reduces the need to continuously (periodically) switch the communication/control structure of distributed control schemes like coalition MPC [9].
@@ VI Conclusion
Some practical issues arise when implementing spectral clustering in large-scale power networks. Firstly, notice that the optimization problem (4) does not explicitly enforce adjacency of the nodes within the resulting clusters. This requirement, commonly referred to as Area Cohesiveness [39], cannot be guaranteed using the framework of spectral clustering. As a consequence, for large values of $k$, some clusters can become disconnected, meaning that the subgraph induced on $\mathcal{G}$ by the corresponding characteristic vector can be disconnected. [40] proposes an algorithm for refining node assignment to ensure that the $k$ clusters obtained using spectral clustering are connected. In future works we intend to address these challenges.
VII Acknowledgment
The authors would like to thank Prof. David Bindel for the helpful conversation on perturbation theory during the finalization of this work.
Appendix A Linearized Dynamics
Given the non-uniform Kuramoto oscillator model
$D_{i}\dot{\delta_{i}}=\omega_{i}-\sum_{j=1}^{n}a_{ij}\sin(\delta_{i}-\delta_{j}),$
where $a_{ij}=\Im(Y_{ij})|V_{i}||V_{j}|$. If we assume the bus voltage magnitudes are constant at $|V_{i}^{*}|$ (not necessarily $1.0$ p.u.), the linearized dynamics of the coupled oscillator model around a synchronous solution $\delta^{*}$ can be obtained as follows.
Let $\Delta\delta_{i}$ and $\Delta\delta_{j}$ be the deviations of the phase angles from the equilibrium so that $\delta_{i}=\delta_{i}^{*}+\Delta\delta_{i}$ and $\delta_{j}=\delta_{j}^{*}+\Delta\delta_{j}$. The first-order Taylor approximation of $\sin(\delta_{i}-\delta_{j})$ around $\delta_{i}^{*}-\delta_{j}^{*}$ is given by
$\sin(\delta_{i}-\delta_{j})\approx\sin(\delta_{i}^{*}-\delta_{j}^{*})+\cos(\delta_{i}^{*}-\delta_{j}^{*})(\Delta\delta_{i}-\Delta\delta_{j}).$
The linearized angle dynamics around $\delta^{*}$ becomes
$D_{i}\dot{\delta_{i}}$ $=D_{i}(\dot{\delta^{*}_{i}}+\dot{\Delta\delta_{i}})=D_{i}\dot{\Delta\delta}_{i}$
$=\omega_{i}-\sum\limits_{j=1}^{n}a_{ij}\big[\sin(\delta_{i}^{*}-\delta_{j}^{*})+\cos(\delta_{i}^{*}-\delta_{j}^{*})(\Delta\delta_{i}-\Delta\delta_{j})\big]$
$=\underbrace{\omega_{i}-\sum\limits_{j=1}^{n}a_{ij}\sin(\delta_{i}^{*}-\delta_{j}^{*})}_{=0}-\sum\limits_{j=1}^{n}a_{ij}\cos(\delta_{i}^{*}-\delta_{j}^{*})(\Delta\delta_{i}-\Delta\delta_{j})$
$=-\sum\limits_{j=1}^{n}a_{ij}\cos(\delta_{i}^{*}-\delta_{j}^{*})(\Delta\delta_{i}-\Delta\delta_{j})$
$=-\sum\limits_{j=1}^{n}a_{ij}\cos(\delta_{i}^{*}-\delta_{j}^{*})\Delta\delta_{i}+\sum\limits_{j=1}^{n}a_{ij}\cos(\delta_{i}^{*}-\delta_{j}^{*})\Delta\delta_{j}$
$=-\sum\limits_{\begin{subarray}{c}{j=1}\\ j\neq i\end{subarray}}^{n}a_{ij}\cos(\delta_{i}^{*}-\delta_{j}^{*})\Delta\delta_{i}+\sum\limits_{\begin{subarray}{c}{j=1}\\ j\neq i\end{subarray}}^{n}a_{ij}\cos(\delta_{i}^{*}-\delta_{j}^{*})\Delta\delta_{j}$
The linearized dynamics for the system can then be written in matrix form as
$D\Delta\dot{\delta}=-\tilde{L}\Delta\delta,$
where $\tilde{L}$ is the laplacian matrix of the graph whose edge weights $\tilde{w}_{ij}=\Im(Y_{ij})|V_{i}||V_{j}|\cos(\delta_{i}^{*}-\delta_{j}^{*})$ and $D$ is the diagonal matrix of damping coefficients.
Appendix B Robustness of the Small Eigenvalues of Linearized System Dynamics Matrix
With a slight abuse of notation, let $L$ denote the laplacian matrix of the dynamic graph, and $D$ denote the diagonal positive definite matrix of damping coefficients.
Given the symmetric pair $(L,D)$, let $\sigma(L,D)=\{\langle\alpha_{1},\beta_{1}\rangle,\langle\alpha_{2},\beta_{2}\rangle,\ldots,\langle\alpha_{n},\beta_{n}\rangle\}$ denote the set of the generalized eigenvalues. $(L,D)$ is definite if

### v2

@@ III-A Linearized Synchronization Dynamics
Since the diagonal matrix of time constants $D$ is non-singular, we can rewrite the linearized synchronization dynamics in state-space form as
@@
where $J$ is the system dynamics matrix. $J$ is similar to the symmetric matrix $D^{-1/2}\tilde{L}D^{-1/2}$ via the transformation matrix $P=D^{-1/2}$, hence they have the same real eigenvalues (the negative of the eigenvalues of the pair $(\tilde{L},D)$) and if $u$ is an eigenvector of $D^{-1/2}\tilde{L}D^{-1/2}$, then $v=Pu=D^{-1/2}u$ is the corresponding eigenvector of $J$ (the generalized eigenvector of the pair $(\tilde{L},D)$). For a connected network with voltage angles $\delta\in\overline{\Delta}$, the zero eigenvalue of $J$ has an algebraic multiplicity of $1$ and the corresponding eigenvector is $\mathbf{1}_{n}$.
For a network of oscillators with identical time constants (i.e., $D=\alpha I$ for some $\alpha>0$), the $i$th eigenvalue of the system dynamics matrix $J$ reduces to the scaled eigenvalues of $\tilde{L}$, that is $\lambda_{i}(J)=-\lambda_{i}(\tilde{L})/\alpha$, while the eigenvectors are exactly the eigenvectors of $\tilde{L}$. In the general case of a network with non-uniform time constants, the Laplacian matrix $\tilde{L}$ and the diagonal matrix of time constants $D$ play a crucial role in determining the evolution of the state variables and the community structure of the network. We show in Section III-B that the eigenvectors corresponding to the smallest eigenvalues of $J$ (also referred to as the inter-area modes of the system) can be used to decompose the network into groups of dynamically coherent nodes.
III-B The Clustering Problem
@@
is the boundary of the set $\mathcal{S}_{i}$ on the graph $\tilde{\mathcal{G}}$, that is, the sum of the edge weights leaving $\mathcal{S}_{i}$ or the total dynamic coupling between $\mathcal{S}_{i}$ and $\mathcal{S}_{i}^{c}$,
@@
is the total damping in the subgraph induced by the set $\mathcal{S}_{i}$. Notice when $D_{i}=1$ for all $i$, $D=I_{n}$, and this term is simply the cardinality of the set $\mathcal{S}_{i}$ denoted by $|\mathcal{S}_{i}|$.
@@
The optimization problem (5) is an integer program known to be computationally challenging for large scale networks [24] and the size of the feasible set, $k^{n}$, grows exponentially with the number of nodes in the network. For any choice of $k$, the solution to (5) corresponds to finding $k$ characteristic vectors in $\{1,0\}^{n}$ with disjoint support [10] that minimizes $\rho(k)$.
@@
$R\left(\tilde{L},D;x_{i}\right)=\frac{x_{i}^{\intercal}\tilde{L}x_{i}}{x_{i}^{\intercal}Dx_{i}}.$ (6)
For any $D\succ 0$ and non-zero vector $x_{i}$ , (6) is well defined; and the relaxation of the integer program (5) is given by
@@
The $k$ vectors minimizing (7) are precisely the eigenvectors corresponding to the $k$ smallest eigenvalues of the pair $(\tilde{L},D)$ [10] with an optimal value of $\psi^{*}(k)=\lambda_{k}(\tilde{L},D)$ – this is a consequence of the generalized Courant-Fischer minimax theorem for the matrix pair $(\tilde{L},D)$ when $\tilde{L}\succeq 0$ and $D\succ 0$ [3, 40]. From (6), we see that the eigenvalues of the pair $(L,D)$ defined as the ratio of the quadratic forms of symmetric psd matrices, are non-negative, therefore the eigenvalues of the system dynamics matrix satisfy $\lambda_{i}(J)\leq 0$ for all $i$.
For a given value of $k$, a power network with a small value of $\psi^{*}(k)$ – the $k$th eigenvalue of the linearized dynamics – implies that the network admits to $k$ loosely coupled partitions with balanced total damping. With $\psi^{*}(k)=0$ if and only if the algebraic multiplicity of the $\lambda_{i}(\tilde{L})=0$ is $k$, that is, if the network has at least $k$ islands.
III-C Generalized Spectral Embedding
Graph embedding is a process of mapping the nodes of a graph to a low dimensional geometric space that preserves the properties of the network. This geometric space, which is generally Euclidean, is such that the distances between coordinates encode the relationships between the nodes in the network. In this section, we use the generalized eigenvectors of the pair $(\tilde{L},D)$, which encodes properties of the linearized dynamics introduced in Section III-B, to construct such an embedding. We then describe a method for obtaining $k$ dynamically coherent clusters (zones) of a power network based on this graph embedding.
@@ III-C Generalized Spectral Embedding
where $\lambda_{i}=\lambda_{i}(\tilde{L},D)\in\mathbb{R}$ and $x_{i}\in\mathbb{R}^{n}$ are the $i$th eigenvalue and right eigenvector, respectively. The generalized eigenvalue problem can be thought of as an ordinary eigenvalue problem over a vector space equipped with a $D$-inner product.
@@
4:Compute $X\in\mathbb{R}^{n\times k}$, the first $k$ eigenvectors of the pair $(\tilde{L},D)$
@@
The number of clusters (the value of $k$) can be determined by the requirements of the underlying control scheme or pre-specified by the system operator. Ideally, $k$ should be such that the first $k$ generalized eigenvalues $\lambda_{1}(\tilde{L},D),\ldots,\lambda_{k}(\tilde{L},D)$ are small while $\lambda_{k+1}(\tilde{L},D)$ is relatively large. In the absence of a prespecified value of $k$, a common heuristic for choosing $k$ is to select $k$ with the largest relative spectral gap, that is
@@
There are several justifications for this choice of $k$. First, a large relative spectral gap ensures that the community structure of the network (the number of clusters) is robust across a wide range of steady-state operating points. Appendix A and B provides an explicit bound on the worst-case change in the eigenvalues of the linearized oscillator model in terms of the spectral radius of the perturbation to the Laplacian of the dynamic graph and a measure of the definiteness of the pair $(\tilde{L},D)$. Secondly, choice of $k$ ensures that the worst of the $k$ clusters( as evaluated using $\phi(\cdot)$) returned by the method is good. We note that the damping coefficients of the oscillators are identical (i.e. $D=\alpha I$, for some $\alpha>0$) the relative spectral gap does not depend on the damping coefficient $\alpha$.
The embedding of the network constructed using Algorithm 1 is unique when the $k$ smallest eigenvalues of the matrix pair $(\tilde{L},D)$ are distinct, since the eigenspaces corresponding to distinct eigenvalues are disjoint subspaces, the $k$ eigenvectors computed at step 4 of Algorithm 1 are therefore unique up to scaling. We note that the dynamic graph $\tilde{\mathcal{G}}$ can be constructed solely from the network parameters and measurements without the need to solve an AC optimal power flow problem. This can significantly reduce the computational complexity of the method to that of computing the eigenvectors of the generalized eigenvalue problem – the most computationally expensive step – which can be computed exactly in polynomial time $\mathcal{O}(n^{3})$ using QZ factorization, or approximately using the inverse power method which converges to a sufficiently good solution in $\mathcal{O}(\log(n))$[10].
@@ IV Results
Figure 1: 22-D Embedding of the IEEE 30-bus test network using the (a) eigenvectors of $\tilde{L}$, (b) the eigenvectors of the pair $(\tilde{L},D)$. The total edge weights cut by the spectral and generalized spectral clustering solutions on $\tilde{\mathcal{G}}$, are 5.04 and 12.55, respectively, and the corresponding total damping in each cluster is 157.89 and 30.01 for spectral clustering, and 91.33 and 96.57 for generalized spectral clustering.
@@ V-A Network Dynamics Simulations
For each $i\in\mathcal{V}$, we perturb the net power injection $\omega_{i}$ and determine the effect of this disturbance on the network by measuring the pair-wise coherence between the angular frequency trajectories $\dot{\delta}_{i}(t)$ and $\dot{\delta}_{j}(t)$ for all $j$ in the synchronously rotating reference frame of frequency $\omega_{\text{sync}}$ defined as
[…]
where $e_{i}(t)=\dot{\delta}_{i}(t)-\omega_{\text{sync}}$, and $\omega_{\text{sync}}=(\sum_{i}\omega_{i})/(\sum_{i}D_{i})$.
@@
V-B Robustness of Clustering Solution
@@
For each load bus in the network, we generate $1,000$ realizations of uniformly distributed load demands as follows
@@
where $\omega_{i}^{0}$ is the nominal demand (in MW) at node $i$, $\zeta_{i}\sim\mathcal{N}(0,\sigma^{2})$ with variance $\sigma^{2}=5$ MW. We then compute the corresponding net-injections of the generator buses, and the voltage magnitudes and angles across the network by solving the ACOPF problem. Randomizing the load demands induces a structured randomization in the realized steady states of the network and the corresponding Laplacian, ensuring the operational feasibility of the resulting states.
Figure 8: Relative spectral gaps for $1,000$ random steady-states of the IEEE 30-bus network.
Across all considered operating points, the largest relative spectral gap was consistently between the $5$th and $6$th generalized eigenvalues. Fig. 8 shows the relative spectral gaps across the realized states. The distribution of the largest relative spectral gap for the network has a mean of $3.1218$ and variance of $0.0213$ (relative variance of $0.007$), confirming that choice of $k=5$ clusters for this network is robust to variability in system operating conditions.
@@
The robustness of the resulting clusters ensures that the identified clusters remain stable under changes in the network’s steady-state operating points, precluding the need to re-solve the clustering problem for slight changes in the network’s loading condition. This reduces the need to periodically switch the communication/control structure of distributed control schemes like coalition MPC [8].
@@ VI Conclusion
Appendix A A Perturbation Bound on Eigenvalues
Given a symmetric pair $(L,D)$, let $\sigma(L,D)=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{n}\}$ denote the set of the generalized eigenvalues. The pair $(L,D)$ is definite if